Please solve this question,
Ben is 3 years older than his sister and a third of the age of his mother.
His father is 4 year older than his mother and four time the age of his sister.
How old is Ben ?
Regards
Raza

Answer by ikleyn(52835)

.
From the condition, you have

B = S + 3,

3B = M,

F = M + 4,

4S = F.


Then moving from the bottom to the top and making substitutions, you get this chain of equalities

4S = F = M + 4 = 3B + 4 = 3*(S+3) + 4.


which gives you 

4S = 3*(S+3) + 4 ====>  4S = 3S + 13  ====>  S= 13.


The sister is 13 years old.  Ben is 13 + 3 = 16 years old.


The mother is  3*16 = 48 years old  and the father is  52 years old.

Solved.